OpenStudy (sloppycanada):

So I'm studying for my wonderful trig final, (yay me!) and I'm going over the whole transformations of a basic trig equation. I don't understand what effect changing the period has on a graph? I'm looking at my textbook and the pictures for 4pi and pi look exactly the same. I know that the higher the frequency (pi) the more waves you get, but it looks the same in my book. No extra lines curves, the two pictures look exactly the same. I emailed my teacher, and she said "they're different, look closely".

OpenStudy (anonymous):

that is why they call it trickonometry

OpenStudy (anonymous):

not sure what "pictures for \(4\pi\) and \(\pi\) look exactly the same" means

OpenStudy (anonymous):

probably just the scale on the x axis is my guess

OpenStudy (sloppycanada):

f(x) = asin(kx - c) + h That's the basic equation, right?

OpenStudy (anonymous):

yeah i guess so now sure why the minus, but if that is what the book has, fine

OpenStudy (sloppycanada):

From what I understand "k" should be the change in period, right? So like 4π or π. But when I look at the graphs sin(4π)x & sin(π)x They look the same to me.

OpenStudy (anonymous):

yeah they do, but look at the scale on the x axis

OpenStudy (sloppycanada):

It actually has + and - but not sure how to make that symbol on the keyboard. same thing for h

OpenStudy (mathmale):

I agree with satellite73 that it's the x-axis scale that differs from graph to graph. The shape of the sine function is always the same, but the period, amplitude, starting point, etc., can all vary. What specific help are you looking for in regard to y = a*sin (kx-c) + h?

OpenStudy (anonymous):

so for example \[\sin(4\pi x)\] has period \[\frac{2\pi}{4\pi}=\frac{1}{2}\]

OpenStudy (sloppycanada):

I'm look at the x-axis, and they both say 180 and 360

OpenStudy (anonymous):

whoa nelly !! really?

OpenStudy (anonymous):

they are graphing with degrees??!!

OpenStudy (sloppycanada):

Yes...?

OpenStudy (anonymous):

holy cow, make them stop

OpenStudy (sloppycanada):

Alright, I'll just go right ahead and email whatever publishing company made my text book. I'll do that right after I get an A on this final.

OpenStudy (anonymous):

sine and cosine are functions, functions of numbers, i.e put in a number, get out a number they only correspond to functions of angles if the angles are measured in radian, not degrees

jimthompson5910 (jim_thompson5910):

@sloppycanada can you show a picture of the graph you're referring to?

OpenStudy (mathmale):

a is the amplitude k is associated with stretching the graph of the sinusoid longer or compressing it so that it is shorter. c has to do with shifting of the graph left or right. h has to do with shifting of the graph up or down.

OpenStudy (anonymous):

yeah what @jim_thompson5910 said, a screen shot would help

OpenStudy (sloppycanada):

Isn't that against the rules? I mean this is a practice worksheet? It's technically cheating, even if I'm not being graded. and I suppose the worksheet is optional.

jimthompson5910 (jim_thompson5910):

As long as it's not a quiz or test or anything, it's fine

OpenStudy (anonymous):

just post it i won't tell btw you did write \(\sin(\pi x)\) right?

OpenStudy (anonymous):

a clear indication that we are working with numbers (or radians if you prefer)

OpenStudy (sloppycanada):

http://prntscr.com/99cg2h

OpenStudy (sloppycanada):

That's the equation I'm working with, and it says to chose a graph, the practice question after it say the same thing but instead of 4π it says π

OpenStudy (sloppycanada):

I've got it narrowed to these two graphs (which thankfully are right next to each other) http://prntscr.com/99cgia

OpenStudy (sloppycanada):

It gives me the same choices for the two questions

OpenStudy (sloppycanada):

I'm pretty sure the changed equation of the 4π is the following - sin(4πx)

OpenStudy (anonymous):

period of \(4\pi\) means it does all it is going to do on an interval of length \(4\pi\) start at 0, go up to one , then back to 0, down to -1, and finally back up to zero in that length

OpenStudy (anonymous):

holy moly you are right!!!!

OpenStudy (anonymous):

it is the first one , but they confused degrees with radians how weird

OpenStudy (mathmale):

In other words, a period of 4Pi involves one complete cycle of either the sine or the cosine function.

OpenStudy (anonymous):

if you call \(-360\) \(-2\pi\) and rename \(360\) as \(2\pi\) then the first graph has period \(4\pi\)

OpenStudy (sloppycanada):

Wait so, the first one is sin(πx)? And then the second one is sin(4πx)?

OpenStudy (anonymous):

although in actuality it is \(720\)

OpenStudy (anonymous):

no, i think there is some confusion here you are not being told the function, just the period

OpenStudy (anonymous):

the one up top, the "longer" one, has period 720

OpenStudy (anonymous):

the one below it, the more compact one, has period 180

OpenStudy (sloppycanada):

Oh so the longer one is the period of 720?

OpenStudy (anonymous):

yes

OpenStudy (sloppycanada):

which means that it'd period is 4π

OpenStudy (sloppycanada):

Okay, see, this is why I like vectors much more.

OpenStudy (anonymous):

they want you to mentally convert \(720^\circ\) to \(4\pi\) because well because they are morons and have no business teaching math

OpenStudy (anonymous):

can you please tell me where this question comes from? what the source is? i

OpenStudy (sloppycanada):

What do you mean?

OpenStudy (anonymous):

is it in a book? by whom is it a worksheet? what is the source?

OpenStudy (sloppycanada):

Ohhh Keystone.

OpenStudy (anonymous):

keystone an on line system? in PA?

OpenStudy (sloppycanada):

Yeah?

OpenStudy (sloppycanada):

Not sure what the relevance is.

OpenStudy (anonymous):

got a reference number for the question ?

OpenStudy (mathmale):

Ladies and gentlemen: Please stick to solving this problem.

OpenStudy (anonymous):

problem is solved unsolved problem is the author of this question, who needs to be contacted

OpenStudy (sloppycanada):

Not sure who made the problem, but I can get the teacher name?

OpenStudy (anonymous):

that would be great, they are usually happy to correct errors, once they are pointed out

OpenStudy (mathmale):

I don't find that the problem posted here is clearly stated. What specificallyl do you want to learn, sloppycanada? Let's leave blame out of this game.

OpenStudy (anonymous):

question was which one had a period of \(4\pi\) it was the "longer " one

OpenStudy (anonymous):

problem was that the graph had the x axis measured in degrees, so saying the period is \(4\pi\) makes no sense

OpenStudy (mathmale):

My understanding is that you're working with y = a*sin(bx + c) + h and want to know how varying each quantity a, b, c and h will affect the shape of the graph. We should be focusing on that.

OpenStudy (sloppycanada):

@mathmale I 've learned what I wanted to do, and solved one of those practice problems on this worksheet.

OpenStudy (mathmale):

If y = a* sin (bx + c), the period is given by the algebraic expression 2Pi/b.

OpenStudy (mathmale):

Good! Further questions?

OpenStudy (sloppycanada):

I wanted to be able to tell the difference between π and 4π or something similar and how that changed the graph. I'm good for now. My brain hurts. I think I'll move to my happy vectors.

OpenStudy (anonymous):

@sloppycanada make sure you do not confuse having a period of \(4\pi\) with the function \[\sin(4\pi x)\] they are not the same thing at all

OpenStudy (mathmale):

If you begin with y=sin x, the period is 2Pi/1, or 2Pi; if you begin with y=sin 4x, the period is 2Pi/(4), or Pi/2.

OpenStudy (sloppycanada):

They're not the same? Then why the heck do I have this stupid "k" in this equation? I thought I had to include that so it made it clear that it was not 4*x

OpenStudy (mathmale):

sloppycanada: Could you possibly share a diagram or drawing? I still find your "question" to be very vague and imprecise. Also, I explained earlier that the quantity k affects the stretching or the compression of the graph of the sine function. the period of sin 2x is half the period of sin x. the period of sin x/2 is twice the period of sin x. Getting to the point where y ou can ask precise questions using the correct math vocabulary will inevitably improve your understanding of trig.

OpenStudy (sloppycanada):

Okay so I suppose it all boils down to the question of - say I have to write a trig equation where they gave me an amplitude of 2, a vertical shift of -4, and a period of 3. How would I write this? Where would I put the period of 3?

OpenStudy (mathmale):

You are GIVEN the period, 3.

OpenStudy (sloppycanada):

So I don't write it the equation?

OpenStudy (mathmale):

start with the graph of y = sin x (or cos x, if that's the case). You need to name the function. y = sin x is the most basic form of this function. If the period is 3, recall that period is given by 2Pi/b. Thus, equate 3 to 2Pi/b and solve for b. If theres's a vertical shift of -4, move the whole graph downward 4 units.

OpenStudy (mathmale):

If the amplitude is to be 2, then draw your sine curve so that it begins at (0,0) and traces a curve as high as 2 units above the x-axis and as low as 2 units beneath the x-axis.

OpenStudy (sloppycanada):

b = 2π/3

OpenStudy (mathmale):

very good. That's the correct b (or k) value when the period is 3. Good!

OpenStudy (sloppycanada):

Yes, but if I had to write the equation of it, 2sin((2π/3)x)) -4?

OpenStudy (mathmale):

Start with y=sin x. If the amplitude is 2, modify that equation to read y=2sin x

OpenStudy (mathmale):

Yes. Good. Must label your expression with "y=" Nice work! y = 2 sin ( (2Pi/3)x ) - 4.

OpenStudy (sloppycanada):

Ohhhhhh so I have to actually change the period of 3 into the 2pi/b thing. got it. Thanks so much for your help.

OpenStudy (mathmale):

Not precisely. Better: If I'm given the fact that the period is 3, I must use the formula for period, 2Pi/b, equating it to the given period (3), and finally solving for b. if b is 2Pi/3, then I write the equation as y = sin (2Pi/3)x.

OpenStudy (mathmale):

If the period is 4, let 4 = 2pi/b and solve for b. And so on. You're welcome! Good luck!! Bye.

OpenStudy (sloppycanada):

So no matter what, I can't just put the period that they give me in, I have to equate it to 2pi/b. I think that's why I've been having problems.

OpenStudy (mathmale):

I'm glad you've made this discovery.